if a,b and c are in arithmetic progression then prove that 2b=a+c?
 

Grade:11

2 Answers

rajesh
8 years ago
since b=a+c/2
therefore 2b=a+c
 
-------------------------------------------------------------------------------------------------------------------------
Vijay Mukati
8 years ago
Dear Student,

If a, b and c are in AP, this means that the comman difference between these numbers will be equal.
i.e. b – a = c – b or 2b = a+c.

Thanks.